# Evidence of Energy Injection in the Short and Distant GRB 250221A

### v4

@@ 2.4 Data Analysis and Photometry
The afterglow of GRB 250221A is projected close to two galaxies (see Section 3.1). Therefore, we performed image subtraction using the Saccadic Fast Fourier Transform (SFFT) (Hu et al., 2022) software using as templates the third epoch of observations with FORS2 for $r$ and archival images from Pan-STARRS DR1 (Magnier et al., 2020) for $g$ and $i$.
The PSF photometry was performed using a custom pipeline and the calibration including SExtractor (Bertin & Arnouts, 1996). To minimize systematic errors in the calibration, we calibrated all our photometry using the Pan-STARRS PS1 DR2 Catalog (Magnier et al., 2020).
@@ 3.1 Environment
Table 2: Properties of the host galaxy candidates of GRB 250221A.
Label RA DEC $z$ $m_{r}$ $R_{\textrm{half}}$ [″] $R_{\rm 0}$ [″] Offset $\delta R$ [kpc] $d_{L}$ [Gpc] $d_{A}$ [Gpc] $P_{\rm ch}(<R_{\rm 0})$
G1 59.4625 -15.1331 0.343 20.86 $\pm$ 0.01 0.77 $11±\pm5$ 1.95 1.06 0.009
G2 59.4628 -15.1333 0.768 24.09 $\pm$ 0.08 0.42666There is no information of the $R_{\textrm{half}}$ due to in the LS catalog, this source is classified as a point source. We assume a radial profile and report the FWHM. 0.29 $7_{-7}^{+12}$ 4.91 1.57 0.003
@@ 4.1.2 Hardness ratio
Figure 5: HR versus $T_{90}$ diagram for GRB 250221A (crimson star), plotted alongside a sample of Swift/BAT GRBs. Division between LGRBs (orange) and SGRBs (blue) is done automatically by Swift/BAT based on $T_{90}$ duration of 2 seconds. Hybrid GRB events (black diamonds) are poorly classified using a purely duration based classification scheme, resulting in ambiguity in progenitor assessment. Both GRB 250221A and GRB 170817A (black star) lie in the ambiguous region between the GRB classes in this space, and within the wide spread of the hybrid GRBs.
@@ 4.1.3 Amati Correlation
Figure 6 shows the resulting relations between the isotropic energy and intrinsic peak energy for GRB 250221A. These values are compared to those in the Swift/BAT GRB database. We distinguish between LGRBs (gray points) and SGRBs (blue points). The Amati relation for LGRBs (solid black line) and the $2\sigma$ (95%) confidence region of the Amati relation for GRBs (Amati et al., 2008; Amati & Della Valle, 2013) are also shown. Redshift values are drawn from the Swift/BAT Gamma-Ray Burst Catalogue [^8] https://swift.gsfc.nasa.gov/results/batgrbcat/, supplemented with spectroscopic measurements from Fong et al. (2022). At the redshifts of both candidate host galaxies, GRB 250221A falls outside the 95%-confidence region for LGRBs and aligns more closely with the short GRB population.
@@
Figure 6: Isotropic energy and intrinsic peak energy for GRB 250221A (pink star). Long GRBs ($T_{90}>2$ seconds) are shown in grey and short GRBs ($T_{90}<2$ seconds) are shown in blue. The Amati relation, based on Swift/BAT GRBs with spectroscopic redshifts, is also shown (solid black line, Amati et al., 2008; Amati & Della Valle, 2013). The $2\sigma$ confidence region of the relation is indicated by dashed lines.
@@ 4.2 Afterglow Model
Within the fireball model framework (Rees & Meszaros, 1992; Meszaros & Rees, 1993; Mészáros & Rees, 1997), the afterglow emission originates from external shocks produced by the interaction of the relativistic jet with the surrounding circumstellar medium (Sari & Piran, 1995; Sari, 1997). Here we consider an ultra-relativistic, thin spherical shell, propagating inside an external medium with constant number density $n$. Initially, the shell is coasting with a bulk Lorentz factor (LF) $\Gamma_{0}\gg 1$ and is slowed down by sweeping up the external medium in its path. This interaction gives rise to a double collisionless shock structure. A forward shock propagates ahead of the shell with $\Gamma_{\rm sh}=\sqrt{2}\Gamma$ and shock-heats the swept up material that moves with bulk LF $\Gamma$ behind it. A reverse shock propagates through the shell and extracts its kinetic energy while shock-heating the ejecta. When the mass of the swept up material exceeds $M_{0}/\Gamma_{0}$, (where $M_{0}=E_{\rm k,iso}/(\Gamma_{0}c^{2})$ is the baryon load of the shell, $E_{\rm k,iso}$ is its kinetic energy, and $c$ the speed of light), the shell starts to decelerate. The subsequent dynamical evolution of the forward shock is given by the Blandford & McKee (1976) self-similar solution, with $\Gamma_{\rm sh}\propto R^{-3/2}$. At this moment most of the kinetic energy of the shell is transferred to the shocked swept up material, resulting in a peak in the afterglow light curve and a power-law decline thereafter.
@@
The forward shock accelerates a fraction of the electrons ($0<\xi_{e}<1$; Eichler & Waxman 2005) entering the shock into a power-law energy distribution, such that their comoving number density behind the shock is $n_{e}^{\prime}(\gamma)\propto\gamma^{-p}$ for electrons with $\gamma\geq\gamma_{m}$ (where $\gamma$ is the LF of the electrons and $\gamma_{m}$ is the LF of minimal-energy electrons). The collisionless shock also compresses any pre-existing magnetic fields in the upstream medium and generates new magnetic fields in-situ. The relativistic electrons behind the shock cool by radiating broadband synchrotron radiation, which is described by a power-law that is broken smoothly at multiple characteristic frequencies (Sari et al., 1998). At late times, the afterglow emission is in the slow-cooling regime where the flux density, $F_{\nu}\propto T^{\alpha}\nu^{\beta}$, peaks at $\nu_{m}<\nu_{c}$, where $T$ is the time in the observer frame and $\nu_{m}$ is the peak synchrotron frequency of minimal-energy electrons (Granot & Sari, 2002). At higher frequencies ($\nu>\nu_{m}$) the spectrum breaks at the characteristic cooling-break frequency, $\nu_{c}$, that corresponds to the peak synchrotron frequency of electrons cooling at the dynamical time. At frequencies $\nu_{m}<\nu<\nu_{c}$, the spectral index is $\beta=(1-p)/2$, while for $\nu>\nu_{c}$, it becomes $\beta=-p/2$. We interpret the observed afterglow radiation in this framework and use the known scaling relations (Sari et al., 1998; Granot & Sari, 2002) to constrain physical parameters.
@@
The early ($T<10^{4}$ s) optical and X-ray data can be modelled with simple power-laws having slopes $\alpha_{\mathrm{o}}=-0.93\pm 0.01$ and $\alpha_{\mathrm{x}}=-0.97\pm 0.04$, respectively (see Figure 1). We interpret this emission to come from the external forward shock and the similarity in the temporal decay suggests that both emissions are produced by the same synchrotron power-law segment (which is at $\nu_{m}<\nu<\nu_{c}$). Comparing the early optical decay slope $\alpha_{\mathrm{o}}=-0.93\pm 0.01$ with the expected one in a constant ambient density ($F\propto T^{3(1-p)/4}$), we obtain an electron power-law index ($p=2.24\pm 0.01$) that is consistent with values reported for other GRBs (e.g., Kumar & Zhang, 2015). This result is incompatible with a wind environment, where the expected decay slope ($F\propto T^{(1-3p)/4}$) would imply $p=1.57\pm 0.01$, an unusually low value for GRBs (see, e.g., Zhang & Mészáros, 2004; Chevalier & Li, 2000).
@@
Figure 7 presents the SED for the early emission phase of GRB 250221A, built using the data from Swift/XRT and COLIBRÍ, corresponding to an average photon arrival time of $T=1800$s for the XRT data. The optical flux densities for the early SED were interpolated from the light curve model described in Section 4.2, evaluated at the same epoch ($T=1800$ s). The X-ray data were binned in groups of at most 10 bins with at least a 10 $\sigma$ significance. Two other epochs at $T=9\times 10^{4}$ s and $T=1.8\times 10^{5}$ s are also shown, and were taken directly from the light curve, as X-ray data were scarce at later times, and a time-sliced spectrum was not possible.
@@
The first two epochs are well described by simple power-laws with spectral index $\beta=-0.71\pm 0.03$ and $\beta=-0.85\pm 0.01$, respectively. The third epoch, covering the peak of the rebrightening in optical and the apparent plateau in X-rays, is best described by two power-law segments: the first one going from the radio to optical energies, with a spectral index $\beta=-0.19\pm 0.02$, the second one going from the optical to X-ray energies, with a spectral index $\beta=-0.91\pm 0.04$.
Therefore, in a slow-cooling synchrotron regime, with an electron power-law index of $p=2.24\pm 0.01$, we obtain an optical spectral index of $\beta=(1-p)/2=-0.62\pm 0.01$. This value differs from the spectral fit $\beta=-0.71\pm 0.03$ at $T+1800$ s, that gives an electron power-law index of $p=2.42\pm 0.06$. The flux density for $\nu_{m}<\nu<\nu_{c}$ at time $T=T_{\rm 1\,day}$ day is thus given by [^9] In this work we use the shorthand notation of $Q_{n}=Q/10^{n}$ for $Q$ expressed in cgs units.
$F_{\nu}\simeq 2.4\times 10^{-3}\left(\frac{1+z}{1.768}\right)^{1.31}\,E_{\rm k,iso,52}^{1.31}\,n_{-1}^{0.5}\,\epsilon_{e,-1}^{1.24}\,\epsilon_{B,-3}^{0.81}\,T_{\rm 1\,day}^{-0.93}\,$ (2)
@@
when assuming $\xi_{e}=1$ and $z=0.768$ with the corresponding luminosity distance of $d_{L}=1.52\times 10^{28}$ cm. Comparing this scaling relation with optical observations at $T=1800$ s and $\nu=4.81\times 10^{14}$ Hz, with $F_{\nu}=6.4\times 10^{-2}$ mJy, allows us to constrain the kinetic energy,
@@
in terms of fiducial parameters.
An additional constraint on $E_{\rm k,iso}$ comes from the efficiency of the prompt $\gamma$-ray emission, $\eta_{\gamma}=[1+(E_{\rm k,iso}/E_{\gamma,\rm iso})]^{-1}\gtrsim 15\%$ (e.g. Beniamini et al., 2016), which would limit the blast wave energy to $E_{\rm k,iso}<[(1-\eta_{\gamma})/\eta_{\gamma}]E_{\gamma,\rm iso}=6.24\times 10^{51}$ erg for if this level of efficiency was realized in this GRB.
Figure 8: (Left) Afterglow model light curve fit to observations using the model from Gill & Granot (2018) that features a blast wave from a spherical shell moving in a ($k=0$) constant density ISM. The 100 thin gray light curves are obtained from sampling the posterior distribution of model parameters, and the thick black and orange light curves are shown for model parameters that best fit the observations. The red vertical line shows the $T_{90}$ duration of the prompt GRB. (Right) Model parameter posterior distributions obtained from an MCMC light curve fit.
@@
assuming that $\Gamma_{0}=100$, which is the minimum LF needed for an on-axis burst to overcome the compactness problem and emit a GRB. SGRBs are typically found with significant offsets from the centers of their respective host galaxies, with 50 percent showing an offset in excess of 5 kpc (e.g. Fong & Berger, 2013; O’Connor et al., 2022). The ISM density is then expected to be low with $10^{-3}\lesssim n({\rm cm}^{-3})\lesssim 1$ (Berger, 2014; O’Connor et al., 2020).
The optical band is above the peak synchrotron frequency of the minimal energy electrons, which is given by Granot & Sari (2002) as
@@
The available data is not able to put strong constraints on the model parameters. To understand the spread in the values of the different parameters, we perform a light curve fit to the optical ($\nu_{O}=3.93\times 10^{14}$ Hz) and X-ray ($\nu_{X}=2$ keV) data using Markov-chain Monte-Carlo (MCMC) methods and a model of a spherical blast wave propagating in a constant density ISM ($k=0$) from Gill & Granot (2018). The fit is obtained using the data before the re-brightening episode at $T\lesssim 3\times 10^{4}$ s and the results are shown in Figure 8.

### v5

@@ 2.4 Data Analysis and Photometry
The afterglow of GRB 250221A is projected close to two galaxies (see Section 3.1). Therefore, we performed image subtraction using the Saccadic Fast Fourier Transform (SFFT) (Hu et al., 2022) software using as templates the third epoch of observations with FORS2 for $r$ and archival images from Pan-STARRS DR1 (Magnier et al., 2020) for $g$ and $i$ (See Figure 2, panel c).
The PSF photometry was performed using a custom pipeline and the calibration including SExtractor (Bertin and Arnouts, 1996). To minimize systematic errors in the calibration, we calibrated all our photometry using the Pan-STARRS PS1 DR2 Catalog (Magnier et al., 2020).
@@ 3.1 Environment
Table 2: Properties of the galaxies in the vicinity of GRB 250221A. We identify G2 as the true host galaxy of GRB 250221A; see Section 3.2.
Label RA DEC $z$ $m_{r}$ $R_{\textrm{half}}$ [″] $R_{\rm 0}$ [″] Offset $\delta R$ [kpc] $d_{L}$ [Gpc] $d_{A}$ [Gpc]
G1 59.4625 -15.1331 0.343 20.86 $\pm$ 0.01 0.77 $11±\pm5$ 1.95 1.06
G2 59.4628 -15.1333 0.768 24.09 $\pm$ 0.08 0.42666There is no information of the $R_{\textrm{half}}$ due to in the LS catalog, this source is classified as a point source. We assume a radial profile and report the FWHM. 0.29 $7_{-7}^{+12}$ 4.91 1.57
@@ 3.2 Determination of the redshift and identification of the host galaxy
Moreover, considering this value of the redshift, this corresponds to a poorly constrained projected offset of <7<7 kpc (see Table 2) which is consistent with typical offsets observed for short GRBs (e.g., Fong et al., 2022; Nugent et al., 2022; O’Connor et al., 2022).
@@ 4.1.2 Hardness ratio
Figure 5: HR versus $T_{90}$ diagram for GRB 250221A (crimson star), plotted alongside a sample of Swift/BAT GRBs (Lien et al., 2016). Division between LGRBs (orange) and SGRBs (blue) is done automatically by Swift/BAT based on $T_{90}$ duration of 2 seconds. Hybrid GRB events (black diamonds) are poorly classified using a purely duration based classification scheme, resulting in ambiguity in progenitor assessment. Both GRB 250221A and GRB 170817A (black star) lie in the ambiguous region between the GRB classes in this space, and within the wide spread of the hybrid GRBs.
@@ 4.1.3 Amati Correlation
Based on the time-averaged spectral analysis performed with the batgrbproduct tool over the interval [-0.220, 1.844] s, the cutoff power-law peak energy in the time-averaged spectral analysis from Swift [^8] https://swift.gsfc.nasa.gov/results/batgrbcat/BAT_refined_circular/1290305/, $E_{\mathrm{p}}=165.2$ keV. Given the limited sensitivity of BAT above $\sim$100–150 keV, the reported $E_{\rm p}$ should be interpreted as a lower limit, indicating that the burst is spectrally hard.
@@
Figure 6 shows the resulting relations between the isotropic energy and intrinsic peak energy for GRB 250221A. These values are compared to those in the Swift/BAT GRB database. We distinguish between LGRBs (gray points) and SGRBs (blue points). The Amati relation for LGRBs (solid black line) and the $2\sigma$ (95%) confidence region of the Amati relation for GRBs (Amati et al., 2008; Amati and Della Valle, 2013) are also shown. Redshift values are drawn from the Swift/BAT Gamma-Ray Burst Catalogue [^9] https://swift.gsfc.nasa.gov/results/batgrbcat/, supplemented with spectroscopic measurements from Fong et al. (2022). Comparing the properties of our burst, we use the value of $z=0.768$ (see Section 3.2) and the values reported by Konus-Wind (Frederiks et al., 2025). Figure 6 shows that GRB 250221A falls outside the 95%-confidence region for LGRBs and aligns more closely with the short GRB population.
@@
Figure 6: Isotropic energy and intrinsic peak energy for GRB 250221A using the values reported by Konus-Wind (Frederiks et al., 2025) (pink star). Long GRBs ($T_{90}>2$ seconds) are shown in grey and short GRBs ($T_{90}<2$ seconds) are shown in blue. The Amati relation, based on Swift/BAT GRBs with spectroscopic redshifts, is also shown (solid black line, Amati et al., 2008; Amati and Della Valle, 2013). The $2\sigma$ confidence region of the relation is indicated by dashed lines.
@@ 4.2 Afterglow Model
Within the fireball model framework (Rees and Meszaros, 1992; Meszaros and Rees, 1993; Mészáros and Rees, 1997), the afterglow emission originates from external shocks produced by the interaction of the relativistic jet with the surrounding circumstellar medium (Sari and Piran, 1995; Sari, 1997). Here we consider an ultra-relativistic, thin spherical shell, propagating inside an external medium with constant number density $n$. Initially, the shell is coasting with a bulk Lorentz factor (LF) $\Gamma_{0}\gg 1$ and is slowed down by sweeping up the external medium in its path. This interaction gives rise to a double collisionless shock structure. A forward shock propagates ahead of the shell with $\Gamma_{\rm sh}=\sqrt{2}\Gamma$ and shock-heats the swept up material that moves with bulk LF $\Gamma$ behind it. A reverse shock propagates through the shell and extracts its kinetic energy while shock-heating the ejecta. When the mass of the swept up material exceeds $M_{0}/\Gamma_{0}$, (where $M_{0}=E_{\rm k,iso}/(\Gamma_{0}c^{2})$ is the baryon load of the shell, $E_{\rm k,iso}$ is its kinetic energy, and $c$ the speed of light), the shell starts to decelerate. The subsequent dynamical evolution of the forward shock is given by the Blandford and McKee (1976) self-similar solution, with $\Gamma_{\rm sh}\propto R^{-3/2}$. At this moment most of the kinetic energy of the shell is transferred to the shocked swept up material, resulting in a peak in the afterglow light curve and a power-law decline thereafter.
@@
The forward shock accelerates a fraction of the electrons ($0<\xi_{e}<1$; Eichler and Waxman 2005) entering the shock into a power-law energy distribution, such that their comoving number density behind the shock is $n_{e}^{\prime}(\gamma)\propto\gamma^{-p}$ for electrons with $\gamma\geq\gamma_{m}$ (where $\gamma$ is the LF of the electrons and $\gamma_{m}$ is the LF of minimal-energy electrons). The collisionless shock also compresses any pre-existing magnetic fields in the upstream medium and generates new magnetic fields in-situ. The relativistic electrons behind the shock cool by radiating broadband synchrotron radiation, which is described by a power-law that is broken smoothly at multiple characteristic frequencies (Sari et al., 1998). At late times, the afterglow emission is in the slow-cooling regime where the flux density, $F_{\nu}\propto T^{\alpha}\nu^{\beta}$, peaks at $\nu_{m}<\nu_{c}$, where $T$ is the time in the observer frame and $\nu_{m}$ is the peak synchrotron frequency of minimal-energy electrons (Granot and Sari, 2002). At higher frequencies ($\nu>\nu_{m}$) the spectrum breaks at the characteristic cooling-break frequency, $\nu_{c}$, that corresponds to the peak synchrotron frequency of electrons cooling at the dynamical time. At frequencies $\nu_{m}<\nu<\nu_{c}$, the spectral index is $\beta=(1-p)/2$, while for $\nu>\nu_{c}$, it becomes $\beta=-p/2$. We interpret the observed afterglow radiation in this framework and use the known scaling relations (Sari et al., 1998; Granot and Sari, 2002) to constrain physical parameters.
@@
The early ($T<3\times 10^{4}$ s) optical and X-ray data can be modelled with simple power-laws having slopes $\alpha_{\mathrm{o}}=-0.93\pm 0.01$ and $\alpha_{\mathrm{x}}=-0.97\pm 0.04$, respectively (see Figure 1). We interpret this emission to come from the external forward shock and the similarity in the temporal decay suggests that both emissions are produced by the same synchrotron power-law segment (which is at $\nu_{m}<\nu<\nu_{c}$). Comparing the early optical decay slope $\alpha_{\mathrm{o}}=-0.93\pm 0.01$ with the expected one in a constant ambient density ($F\propto T^{3(1-p)/4}$), we obtain an electron power-law index ($p=2.24\pm 0.01$) that is consistent with values reported for other GRBs (e.g., Kumar and Zhang, 2015). This result is incompatible with a wind environment, where the expected decay slope ($F\propto T^{(1-3p)/4}$) would imply $p=1.57\pm 0.01$, an unusually low value for GRBs (see, e.g., Zhang and Mészáros, 2004; Chevalier and Li, 2000).
@@
Figure 7 presents the SED for the early emission phase of GRB 250221A, built using the data from Swift/XRT and COLIBRÍ, corresponding to an average photon arrival time of $T=1800$s for the XRT data. The optical flux densities for the early SED were interpolated from the light curve model described in the previous paragraph, evaluated at the same epoch ($T=1800$ s). The X-ray data were binned in groups of at most 10 bins with at least a 10 $\sigma$ significance. Two other epochs at $T=9\times 10^{4}$ s and $T=1.8\times 10^{5}$ s are also shown, and were taken directly from the light curve, as X-ray data were scarce at later times, and a time-sliced spectrum was not possible.
@@
The first two epochs are well described by simple power-laws with spectral index $\beta=-0.71\pm 0.03$ and $\beta=-0.85\pm 0.01$, respectively. The third epoch, covering the peak of the rebrightening in optical and the apparent plateau in X-rays, is best described by two power-law segments: the first one going from the radio to optical energies, with a spectral index $\beta=-0.19\pm 0.02$, the second one going from the optical to X-ray energies, with a spectral index $\beta=-0.91\pm 0.04$. Given the limited number of data used to infer the radio-to-optical spectral slope, the exact position of the spectral break is uncertain and the low-energy spectral slope (-0.19) is not well constrained and should be interpreted as a lower limit.
Therefore, in a slow-cooling synchrotron regime, with an electron power-law index of $p=2.24\pm 0.01$, we obtain an optical spectral index of $\beta=(1-p)/2=-0.62\pm 0.01$. This value differs from the spectral fit $\beta=-0.71\pm 0.03$ at $T+1800$ s, that gives an electron power-law index of $p=2.42\pm 0.06$. The flux density for $\nu_{m}<\nu<\nu_{c}$ at time $T=1800~s(T_{3.26}$ s [^10] In this work we use the shorthand notation of $Q_{n}=Q/10^{n}$ for $Q$ expressed in cgs units.) is thus given by (Granot and Sari, 2002)
$F_{\nu}\simeq 0.087\left(\frac{1+z}{1.768}\right)^{1.31}\,E_{\rm k,iso,52}^{1.31}\,n_{-1}^{0.5}\,\epsilon_{e,-1}^{1.24}\,\epsilon_{B,-3}^{0.81}\,T_{3.26}^{-0.93}\,$ (2)
@@
when assuming fiducial values for the model parameters and for $\xi_{e}=1$ and $z=0.768$ with the corresponding luminosity distance of $d_{L}=1.52\times 10^{28}$ cm. Comparing this scaling relation with optical observations at $T=1800$ s and $\nu=4.81\times 10^{14}$ Hz, with $F_{\nu}=6.4\times 10^{-2}$ mJy, allows us to constrain the kinetic energy,
@@
in terms of fiducial parameters whose values are not yet determined from other constraints.
An additional constraint on $E_{\rm k,iso}$ comes from the efficiency of the prompt $\gamma$-ray emission, $\eta_{\gamma}=[1+(E_{\rm k,iso}/E_{\gamma,\rm iso})]^{-1}\gtrsim 15\%$ (e.g. Beniamini et al., 2016), which would limit the blast wave energy to $E_{\rm k,iso}<[(1-\eta_{\gamma})/\eta_{\gamma}]E_{\gamma,\rm iso}=6.24\times 10^{51}$ erg if this level of efficiency was realized in this GRB. [^11] However, we do not apply this efficiency constraint (which is not robust enough and can vary among different GRBs, especially those showing a rebrightening).
Figure 8: (Left) Afterglow model light curve fit to observations using the model from Gill and Granot (2018) that features a blast wave from a spherical shell moving in a ($k=0$) constant density ISM. The 100 thin gray light curves are obtained from sampling the posterior distribution of model parameters, and the thick black and orange light curves are shown for model parameters that best fit the observations. The red vertical line shows the $T_{90}$ duration of the prompt GRB. (Right) Model parameter posterior distributions obtained from an MCMC light curve fit.
@@
assuming that $\Gamma_{0}=100\,\Gamma_{0,2}$, which is the minimum LF needed for an on-axis burst to overcome the compactness problem and emit a GRB. SGRBs are typically found with significant offsets from the centers of their respective host galaxies, with 50 percent showing an offset in excess of 5 kpc (e.g. Fong and Berger, 2013; O’Connor et al., 2022). The ISM density is then expected to be low with $10^{-3}\lesssim n({\rm cm}^{-3})\lesssim 1$ (Berger, 2014; O’Connor et al., 2020). In fact, O’Connor et al. (2020) find that the central 60 per cent of SGRBs in their sample have ISM densities in the range $10^{-4}\lesssim n_{\rm ISM}/{\rm cm^{-3}}\lesssim 10^{-1.5}$ and only a minute fraction have densities above $1\,{\rm cm^{-3}}$. Therefore, the minimum ISM density found for this source (see MCMC fit below) represents the extreme and may even disfavour a compact binary merger scenario.
The optical band is above the peak synchrotron frequency of the minimal energy electrons, which is given by Granot and Sari (2002) as
@@
When assuming the minimum value of $\epsilon_{B}$ we constrain the minimum ISM density using Eq.(5) to $n_{\rm min}\simeq 2.3\times 10^{3}\Gamma_{0,2}^{-8.01}\,{\rm cm^{-3}}$. This density is extremely high and never realized for GRBs. It can be reduced to a more acceptable value by raising $\Gamma_{0}$ to a larger value, e.g. $\Gamma_{0}=400$ (as obtained below from our MCMC fit), which yields $n_{\min}\simeq 0.034\Gamma_{0,2.6}^{-8.01}\,{\rm cm^{-3}}$. We then use the minimum ISM density to obtain a constraint on the isotropic-equivalent kinetic energy of the blast wave from above using Eq.(3), $E_{\rm k,iso,max}\simeq 5.5\times 10^{54}\Gamma_{0,2.6}^{-0.02}$ erg (Here and above we have ignored the dependence on $\epsilon_{e}$ as it has a very small power). Alternatively, we consider the maximum value of $\epsilon_{B}$ and find $n_{\min}\simeq 3.5\epsilon_{e,-1}^{-1.25}\Gamma_{0,2}^{-8.3}\,{\rm cm^{-3}}$, which further yields $E_{\rm k,iso,max}\simeq 8.9\times 10^{51}\epsilon_{e,-1}^{-1.24}\Gamma_{0,2}^{-0.34}$ erg. The available data is not able to put strong constraints on the model parameters. To understand the spread in the values of the different parameters, we perform a light curve fit to the optical ($\nu_{O}=3.93\times 10^{14}$ Hz) and X-ray ($\nu_{X}=2$ keV) data using Markov-chain Monte-Carlo (MCMC) methods and a model of a spherical blast wave propagating in a constant density ISM ($k=0$) from Gill and Granot (2018). The fit is obtained using the data before the re-brightening episode at $T\lesssim 3\times 10^{4}$ s and the results are shown in Figure 8.